# An extrapolation-driven network architecture for physics-informed deep learning

### v3

@@ Abstract
Current PINN versions with sequential learning strategies have some weaknesses, such as the failure to reproduce the previous training results by a single network, the difficulty to strictly ensure continuity and smoothness at the time interval nodes by multiple networks, and the increase in complexity and computational overhead. To compensate for these shortcomings, we first investigate the extrapolation capability of the PINN method for time-dependent PDEs. Taking advantage of this extrapolation property, we can generalize the training result obtained in a specific time subinterval to larger intervals by adding a correction term to the network parameters of the subinterval. The correction term is determined by further training with the sample points in the added subinterval. Secondly, by designing an extrapolation control function with special characteristics and combining it with a correction term, we construct a new neural network architecture whose network parameters are coupled with the time variable, which we call the extrapolation-driven network architecture. Based on this architecture, using a single neural network, we can obtain the overall PINN solution of the whole domain with the following two characteristics: (1) it completely inherits the local solution of the interval obtained from the previous training, (2) at the interval node, it strictly maintains the continuity and smoothness that the true solution has. The extrapolation-driven network architecture allows us to divide a large time domain into multiple subintervals and solve the time-dependent PDEs one by one in a chronological order. This training scheme respects the causality principle and effectively overcomes the difficulties of the conventional PINN method in solving the evolution equation on a large time domain. Numerical experiments verify the performance of our proposed method. The data and code accompanying this paper are available at https://github.com/wangyong1301108/E-DNN.
@@ 1 Introduction
We focus here on solving time-dependent PDEs by PINNs. In Ref. [13], Penwarden et al. argue that the conventional PINN method faces some difficulties when trained to predict over the entire spatio-temporal domain. They summarize these difficulties into 3 types: (1) Zero solution, where PINNs tend to converge to a zero solution when the zero solution could minimize residual loss; (2) No propagation, where PINNs fail to propagate any information due to an insufficient number of residual points, resulting in invalid predictions;(3) Incorrect propagation, where PINNs quickly converge to an incorrect solution, perhaps because the optimizer falls into a local minimum. To address these difficulties, researchers have proposed many improved versions of the conventional PINN method, including domain decomposition strategies [14, 15, 16], sequential learning strategies [13, 17, 18, 19], and causal weights [20]. While the method in [20] has been successful in simulating complex systems, its training time is very long. As a positive development, Ref. [21] mitigates the problem by modifying the causality parameter. Additionally, sequential learning strategies are often preferred in applications [22, 23] because they naturally respect the principle of causality [20] and have a relatively higher computational efficiency.
@@
The extrapolation has been concerned in multilayer perceptrons (MLPs)  [27, 28] and PINNs [29, 30, 31]. Prior work demonstrates that MLPs usually succeed in interpolation, but fail to extrapolate well [32, 33]. However, as mentioned in Refs. [29, 30], PINNs are capable of interpolation and extrapolation by using physical information as soft penalty constraints. Ref. [31] illustrates that PINNs have credible interpolation and extrapolation properties when applied to velocity models containing gaps. Designing appropriate neural network architectures based on the extrapolation property of PINNs is a feasible way to address the above issues.
@@ 4.1 Allen-Cahn equation with highly nonlinear source term
Table 4.2 lists the results of various methods, including the $L_{2}$ relative error and training time. Considering that the results in the literature generally use the float32 data type, we also give the computational result using the float32 data type for this test. Compared to the other methods proposed in the recent literature, our methods provide more accurate predictions and show better efficiency. Compared to the sE-PINN method, the aE-PINN method introduces an additional trainable parameter ($T_{f}$), which slightly increases the training time.
@@ 4.2 Convection equation with large convection coefficient
Table 4.4 lists the results of different PINN versions. As can be seen, our approaches have a better performance for this PDE as well.
@@ 5 Conclusion
To address the limitations of sequential learning strategies, we systematically analyzed the extrapolation property of the PINN solution, then motivated by the extrapolation capability, we propose a new neural network architecture called E-DNN. The new PINN approaches derived from the E-DNN architecture can perfectly address the aforementioned problems. Moreover, our approaches are relatively computationally efficient and have inherent scalability. With the proposed approaches, using only a single neural network, we can divide the large time domain into several subintervals and then obtain the local PINN solution of each of the subintervals one by one, and the training process is independent among them; especially, the overall solution of the whole time domain combined by these local PINN solutions is seamless, it strictly preserves the continuity and smoothness at the interval nodes. In addition, our approach has good scalability and is naturally suited for high-dimensional problems. However, it also has some limitations. One of the potential problems is that in the process of one-by-one training, the error of the previous time interval could be deposited in the next time interval, resulting in a certain amount of error accumulation. How to reduce or eliminate the possible error accumulation is our future work.
@@ Acknowledgements
For Eq. (2.3), Figure A.1 shows the extrapolation performance of $\frac{T}{2}$ obtained with different network sizes for a fixed number of training points. Figure A.2 gives the results of different numbers of training points for a fixed neural network size. Figures A.3 and A.4 give the similar results for Eq. (2.4) with $\beta=40$.

### v4

@@ Abstract
Current physics-informed neural network (PINN) implementations with sequential learning strategies often experience some weaknesses, such as the failure to reproduce the previous training results when using a single network, the difficulty to strictly ensure continuity and smoothness at the time interval nodes when using multiple networks, and the increase in complexity and computational overhead. To overcome these shortcomings, we first investigate the extrapolation capability of the PINN method for time-dependent PDEs. Taking advantage of this extrapolation property, we generalize the training result obtained in a specific time subinterval to larger intervals by adding a correction term to the network parameters of the subinterval. The correction term is determined by further training with the sample points in the added subinterval. Secondly, by designing an extrapolation control function with special characteristics and combining it with a correction term, we construct a new neural network architecture whose network parameters are coupled with the time variable, which we call the extrapolation-driven network architecture. Based on this architecture, using a single neural network, we can obtain the overall PINN solution of the whole domain with the following two characteristics: (1) it completely inherits the local solution of the interval obtained from the previous training, (2) at the interval node, it strictly maintains the continuity and smoothness that the true solution has. The extrapolation-driven network architecture allows us to divide a large time domain into multiple subintervals and solve the time-dependent PDEs one by one in a chronological order. This training scheme respects the causality principle and effectively overcomes the difficulties of the conventional PINN method in solving the evolution equation on a large time domain. Numerical experiments verify the performance of our method. The data and code accompanying this paper are available at https://github.com/wangyong1301108/E-DNN.
@@ 1 Introduction
We focus here on solving time-dependent PDEs by PINNs. In Ref. [13], Penwarden et al. argue that the conventional PINN method faces some difficulties when trained to predict over the entire spatio-temporal domain.
They summarize these difficulties into 3 types: (1) Zero solution, where PINNs tend to converge to a zero solution when the zero solution could minimize residual loss; (2) No propagation, where PINNs fail to propagate any information due to an insufficient number of residual points, resulting in invalid predictions;(3) Incorrect propagation, where PINNs quickly converge to an incorrect/trivial solution, perhaps because the optimizer falls into a local minimum corresponding to a trivial soluton. To address these difficulties, researchers have proposed many improved versions of the conventional PINN method, including domain decomposition strategies [14, 15, 16], sequential learning strategies [13, 17, 18, 19], and causal weights [20]. While the method in [20] has been successful in simulating complex systems, its training time is very long. As a positive development, Ref. [21] mitigates the problem by modifying the causality parameter. Additionally, sequential learning strategies are often preferred in applications [22, 23] because they naturally respect the principle of causality [20] and have a relatively higher computational efficiency.
The extrapolation properties of multilayer perceptrons (MLPs)  [27, 28] and PINNs [29, 30, 31] have been widely studied. Prior work demonstrates that MLPs usually succeed in interpolation, but fail to extrapolate well [32, 33]. However, as mentioned in Refs. [29, 30], PINNs are capable of interpolation and extrapolation by using physical information as soft penalty constraints. Ref. [31] illustrates that PINNs have credible interpolation and extrapolation properties when applied to velocity models containing gaps. Designing appropriate neural network architectures based on the extrapolation property of PINNs is a feasible way to address the above issues.
@@ 4.1 Allen-Cahn equation with highly nonlinear source term
Table 4.2 lists the accuracy (given by the $L_{2}$ relative error) and training time for the various methods. Considering that the float32 data type is often used in the literature, we also use the float32 data type for this test. Compared to the other methods proposed in recent years, our methods provide more accurate predictions and show better efficiency. Compared to the sE-PINN method, the aE-PINN method introduces an additional trainable parameter ($T_{f}$), which slightly increases the training time.
@@ 4.2 Convection equation with large convection coefficient
Table 4.4 lists the results of different PINN versions. As we can see, our approaches have a better performance for this PDE as well.
@@ 5 Conclusion
To address the limitations of sequential learning strategies, we systematically analyzed the extrapolation property of the PINN solution, then motivated by the extrapolation capability, we propose a new neural network architecture called E-DNN. The new PINN approaches derived from the E-DNN architecture can perfectly address the aforementioned problems. Moreover, our approaches are relatively computationally efficient and have inherent scalability. With the proposed approaches, using only a single neural network, we can divide the large time domain into several subintervals and then obtain the local PINN solution of each of the subintervals one by one, and the training process is independent among them; especially, the overall solution of the whole time domain combined by these local PINN solutions is seamless, it strictly preserves the continuity and smoothness at the interval nodes. In addition, our approach has good scalability and is naturally suited for high-dimensional problems. However, it also has some limitations. One of the potential problems is that in the process of one-by-one training, the error of the previous time interval could be propagated into the next time interval, resulting in a certain amount of error accumulation. How to reduce or eliminate the possible error accumulation is our future work.
@@ Acknowledgements
For Eq. (2.3), Figure A.1 shows the extrapolation performance of $\frac{T}{2}$ obtained with different network sizes for a fixed number of training points. Figure A.2 gives the results of different numbers of training points for a fixed neural network size. Figures A.3 and A.4 display the similar results for Eq. (2.4) with $\beta=40$.